# Wigner distribution, Wigner entropy, and Anomalous Transport of a Generalized Aubry-André model

### v2

@@ IV Quantum transport property
In the previous study, it was proposed that the mobility edge can observed by the dynamical observation Xu et al. (2020) in an one-dimensional incommensurate system with generalized quasidisordered potential, where extended-critical transition can be characterized by the transition from the ballistic transport to the absence of diffusion. In fact, in addition to the ballistic transport and absence of diffusion, there are anomalous transport phenomena in the incommensurate system Purkayastha et al. (2021); Chiaracane et al. (2021); Purkayastha et al. (2018); Lacerda et al. (2021), such as the superdiffusive and subdiffusive transports. Therefore, in this section, we are concerned about whether anomalous transport occurs in the system with the generalized quasidisordered potential similar to that in Ref. Xu et al. (2020) and with the invariable mobility edge.
We analyze the transport behaviors by studying the dynamical evolution of the wave packet. The isolated generalized AA model is initialized with a particle occupying on site $L/2$. Thus, the initial state |$\ket{\psi(\tau=0)}$ is described by the wave function |$\ket{\psi(\tau=0)}=\sum_{n}\phi_{n}(\tau=0)\hat{c}^{{\dagger}}_{n}\ket{0}$ where $\phi_{n}(\tau=0)=\delta_{n,L/2}$ is the Kronecker $\delta$ function. By solving the time-dependent Schrödinger equation, we obtain the evolution of the wave function probability amplitude with time $\phi_{n}(\tau)$, and then the mean squared displacement of wave function is given by Chiaracane et al. (2021)
@@
$\Delta x^{2}(\tau)\sim\tau^{\gamma},$ (6)
where $\gamma=2$ represents the ballistic transition, $\gamma=0$ implies absence of diffusion, $\gamma=1$ denotes diffusion, $10<\gamma<$ 1 shows subdiffusive transport, and $21<\gamma<$ 2 corresponds to superdiffusive transport. In Figs. 5(a)-5(f), we plot the the time evolution of $\Delta x^{2}(\tau)$ under $b=0.4$ and $L=1000$. From the results of $\Delta x^{2}(\tau)$, we conclude the in this one-dimensional incommensurate system with generalized quasidisordered potential and invariant mobility edge, the evolution of wave packet is not only ballistic transport and absence of diffusion, but also anamolous transport phenomena. Specifically, when the potential strength is small, such as $V=0.2t$, the system is in the extended phase. Then we can see from Fig. 5(a) that the evolution of wave packet presents ballistic transport. As $V$ get larger, the system entries the mobility edge regime, there are anomalous transport phenomena. From Figs. 5(b)-5(e), we can see that the evolution curves of $\Delta x^{2}(\tau)$ presents superdiffusion, diffusion, and subdiffusion transport in turn. In the localization regime, such as $V=10t$, $\Delta x^{2}(\tau)$ scales as $\tau^{0}$ (see Fig. 5(f)), meaning the absence of diffusion.
@@ V Summary
In conclusion, we have studied a generalized Aubry-André model with tunable quasidisordered potentials. The model exists an invariable mobility edge that can separate the extended state from the localized state. At the invariable mobility edge, the wave functions are critical. We have proved that one can employ the physical image in phase space, i.e., the Wigner distribution, to distinguish the extended states, critical states, and the localized states. Based on the Wigner distribution, we can further obtain the Wigner entropy. Our numerical results show that it is also feasible to use Wigner entropy to distinguish extended, critical and local states. The critical state has the largest entropy, the extended state the middle, and the local state the smallest. Using this property, we can locate the invariable mobility edge. In addition, by studying the wave packet dynamics of this model, we find that there are not only ballistic transport, diffusion, and absence of diffusion, but also are superdiffusion and subduffusion. These results further expand the extended-localized transition and dynamic properties of generalized quasidisordered systems. In the near future, we hope that these approaches can be generalized to other disordered or quasidisordered systems to distinguish between different forms of phase transitions.
@@ Acknowledgements.
This work is supported by the Natural Science Foundation of Zhejiang Province (Grants No. LQN25A040012) and the start-up fund from Xingzhi college, Zhejiang Normal University, and the National Natural Science Foundation of China (Grants No. 12174346),

### v3

@@ IV Quantum transport property
In the previous study, it was proposed that the mobility edge can observed by the dynamical observation [16] in an one-dimensional incommensurate system with generalized quasidisordered potential, where extended-critical transition can be characterized by the transition from the ballistic transport to the absence of diffusion. In fact, in addition to the ballistic transport and absence of diffusion, there are anomalous transport phenomena in the incommensurate system [30, 31, 32, 33], such as the superdiffusive and subdiffusive transports. Therefore, in this section, we are concerned about whether anomalous transport occurs in the system with the generalized quasidisordered potential and with a mobility edge.
Figure 6: (Color Online) Time evolution of mean squared displacement $\Delta x^{2}(\tau)$ from the system of (1) for $b=0.4$ and $L=1000$. (a) $V=0.1t$; (b) $V=5.2t$; (c) $V=7.3t$; (d) $V=7.5t$; (e) $V=7.6t$; and (f) $V=12t$.
Figure 7: (Color Online) Time evolution of mean squared displacement $\Delta x^{2}(\tau)$ from the system of (5) for $p=1.5$ and $L=500$. (a) $V=0.05t$; (b) $V=0.65t$; (c) $V=0.7t$; (d) $V=0.75t$; (e) $V=0.85t$; and (f) $V=1.2t$.
[…]
@@
$\Delta x^{2}(\tau)\sim\tau^{\beta},$ (7)
where $\beta=2$ represents the ballistic transport, $\beta=0$ implies absence of diffusion, $\beta=1$ denotes diffusion, $0<\beta<1$ shows subdiffusive transport, and $1<\beta<2$ corresponds to superdiffusive transport.
We first examine the wave packet dynamics in the system described by Eq. 1. For parameters $b=0.4$ and $L=1000$, the mean squared displacement $\Delta x^{2}(\tau)$ is plotted for various $V/t$ in Figs. 6(a)-6(f), where the red lines denote the fitting curves using $\log_{10}(\Delta x^{2})=\beta\log_{10}(\tau)+\gamma$. The fitting parameter and associated errors are provided in Appendix A. As shown in Fig. 6(a), ballistic transport dominates for small potential strengths. With increasing $V$, the dynamics transition progressively through superdiffusive, diffusive, and subdiffusive regimes [Figs. 6(b)-(e)], culminating in non-diffusive behavior characterized by sharp fluctuations around the mean [Fig. 6(f)] for sufficiently large $V$.
Similar results are obtained for the Biddle-Sarma model. For $p=1.5$ and $L=500$, $\Delta x^{2}(\tau)$ is shown for different $V$ in Figs. 7(a)-7(e). Ballistic transport prevails at small $V$ [Fig.7(a)], evolving into superdiffusive, diffusive, and subdiffusive transport as $V$ increases [Figs.7(b)–(e)]. At large $V$, $\Delta x^{2}(\tau)$ oscillates near equilibrium, indicative of non-diffusive dynamics [Fig.7(f)]. These findings demonstrate that, in generalized AA models with mobility edges, transport encompasses not only ballistic, diffusive, and non-diffusive regimes but also anomalous superdiffusive and subdiffusive behaviors.
@@ V Summary
In conclusion, we have studied a generalized Aubry-André model with tunable quasidisordered potentials. The model exists an invariable mobility edge that can separate the extended state from the localized state. At the invariable mobility edge, the wave functions are critical. We have proved that one can employ the physical image in phase space, i.e., the Wigner distribution, to distinguish the extended states, critical states, and the localized states. Based on the Wigner distribution, we can further obtain the Wigner entropy. Our numerical results show that it is also feasible to use Wigner entropy to distinguish extended, critical and local states. The critical state has the largest entropy, the extended state the middle, and the local state the smallest. Using this property, we can locate the invariable mobility edge. Meanwhile, this method is again verified in Biddl-Sarma model with exact mobility edge. In addition, by studying the wave packet dynamics of the generalized AA model and the Biddle-Sarma model, we find that there are not only ballistic transport, diffusion, and absence of diffusion, but also are superdiffusion and subduffusion. These results further expand the extended-localized transition and dynamic properties of generalized quasidisordered systems. In the near future, we hope that these approaches can be generalized to other disordered or quasidisordered systems to distinguish between different forms of phase transitions.
@@ Acknowledgements.
We thank Tan Hailin for insightful discussions. This work is supported by the Natural Science Foundation of Zhejiang Province (Grants No. LQN25A040012) and the start-up fund from Xingzhi college, Zhejiang Normal University, and the National Natural Science Foundation of China (Grants No. 12174346),
Appendix A Fitting Parameters
Table 1 lists the fitted coefficients $\beta$ and $\gamma$, along with their standard errors, for the function $\log_{10}[\Delta x^{2}(\tau))]=\beta\log_{10}(\tau)+\gamma$ , where $\beta$ and $\gamma$.
Table 1: Coefficients and standard errors of the fitting function
Fig $\beta/\gamma$ Coefficient Value $\beta/\gamma$ Standard Error
Fig6(a) 1.9878/0.3080 0.0001/0.0001
Fig6(b) 1.9004/-1.9030 0.0009/0.0023
Fig6(c) 1.0174/-1.5884 0.0006/0.0020
Fig6(d) 0.6820/-1.1141 0.0008/0.0027
Fig6(e) 0.2982/-0.1953 0.0010/0.0037
Fig6(f) 0.0029/-1.2738 0.0015/0.0054
Fig7(a) 1.9910/-0.9143 0.0000/0.0001
Fig7(b) 1.5975/-2.3576 0.0008/0.0023
Fig7(c) 1.3402/-2.0962 0.0009/0.0026
Fig7(d) 0.9540/-1.5658 0.0019/0.0045
Fig7(e) 0.5455/-1.2127 0.0004/0.0015
Fig7(f) -0.0000/-1.2232 0.0018/0.0070